OpenStudy (anonymous):

what 2 numbers when multiplied = 52, but when added = -25?

OpenStudy (anonymous):

if you call those numbers x and y, you can create equations: xy=52 x+y=-25 The bottom equation can be solved for x or y x=-25-y And this can be substituted into the first equation. (-25-y)y=52 Solve for y. Once you have y, substitute into either of the two original equations and you'll get x.